### v2

# A Methodological Report on Anomaly Detection on Dynamic Knowledge Graphs

Abstract
In this paper, we explore different approaches to anomaly detection on dynamic knowledge graphs, specifically in a microservices environment for Kubernetes applications. Our approach explores three dynamic knowledge graph representations: sequential data, one-hop graph structure, and two-hop graph structure, with each representation incorporating increasingly complex structural information. Each phase includes different machine learning and deep learning models. We empirically analyse their performance and propose an approach based on ensemble learning of these models. Our approach significantly outperforms the baseline on the ISWC 2024 Dynamic Knowledge Graph Anomaly Detection dataset, providing a robust solution for anomaly detection in dynamic complex data.
Keywords:
Dynamic Knowledge Graph Anomaly Detection.
1 Introduction
With the development of graph machine learning and geometric deep learning [1], research on encoding and prediction has been applied to more domains than ever before. Graphs can simultaneously represent entities through nodes [^1] Also referred to as ”vertices”. and relationships between entities through edges. As a result, graph representations are widely used in domains as diverse as transportation [2], chemistry [3], and social networks [24]. Two areas of relative difficulty in graph learning are heterogeneous graphs and dynamic graphs. Heterogeneous graphs allow for different types of nodes and/or edges in the graph [4]. They can therefore better model complex systems with different types of entities, such as a distributed information system with services, pods and nodes. Dynamic graphs allow the structure and/or attributes of a graph to change over time [5], allowing better modelling of data with both temporal and structural information.
Dynamic Knowledge Graphs (DKG), as an intersection of heterogeneous and dynamic graphs, have attracted increasing attention due to their ability to represent the evolving nature of knowledge over time. By exploiting the ability to capture topological and attribute transformations, DKGs provide valuable insights into the evolution of knowledge in diverse domains such as social media [23], the Internet of Things (IoT) [22, 21], and e-commerce [^2] https://eval.ai/web/challenges/challenge-page/2267/overview.
On the leading Micro-Services Architecture (MSA) platform, such as Kubernetes, there are different types of entities such as compute nodes, deployable units (pods), and services that work together to build applications such as e-commerce platforms. These data are typically stored in DKGs in the TTL format [^3] https://en.wikipedia.org/wiki/Turtle_(syntax). Predictive analysis based on such data is critical to the operation and optimisation of these platforms, particularly for anomaly detection tasks aimed at accurately identifying various types of anomalies.
Despite rapid progress in commercial applications, as a relatively young field, there is still a lack of academic datasets and methodologies. This scarcity can be attributed to several challenges: anomaly events are typically very rare, resulting in an unbalanced data distribution that makes it difficult for models to effectively learn anomaly features and affects their generalisation ability in detecting new anomalies. In addition, models need to capture temporal and structural information and handle heterogeneous nodes/edges, which makes common RNNs and homogeneous GNNs inadequate.
To address these challenges, this paper introduces a novel anomaly detection framework for dynamic knowledge graphs, leveraging both graph structure and temporal dependencies. Our method integrates neighbourhood feature extraction with advanced machine learning techniques, including Support Vector Machine (SVM) [6], Extreme Gradient Boosting (XGB) [8], and Self-Attention mechanism (SA) [13] to enhance the model’s ability to capture complex patterns of anomalies. Our approach achieves significant performance improvements over the baseline in the Anomaly Detection on Dynamic Knowledge Graphs challenge (ADDKG) [^4] https://eval.ai/web/challenges/challenge-page/2267/overview at ISWC 2024, providing a robust solution for anomaly detection in dynamic and complex environments.
The structure of the paper is as follows. Section 2 gives the formulation of the problem and describes related work. Section 3 presents the methodology of the paper. Section 4 describes the experimental part of the work and discusses the results obtained. Section 5 concludes our contribution to the ADDKG challenge and discusses future work.
2 Related Work
Recent studies have shown that the dynamic knowledge graph representation is expressive for anomaly detection in temporally complex systems. In this section, we review the related work in this area. In subsection 2.1, we present the definition and notations of the Dynamic Knowledge Graph. Then, in subsection 2.2, we present common approaches to anomaly detection.
2.1 Dynamic Knowledge Graphs
In static graphs, common representations include edge lists and adjacency matrices. Among them, the edge list is widely used in software development as it requires less memory. An edge list is a collection of edges, each represented as a pair of nodes.
In the realm of dynamic graphs, existing research categorises them into Continuous-Time Dynamic Graphs (CTDG) and Discrete-Time Dynamic Graphs (DTDG). CTDGs represent dynamic graphs using timestamped events, such as $(u,v,t)$, where $u$ and $v$ are nodes and $t$ is the timestamp of the interaction. In contrast, DTDGs represent dynamic graphs as a sequence of static graph snapshots $\{G_{1},G_{2},...,G_{T}\}$ at different time intervals.
Combined with the dynamic graph representations, the Dynamic Knowledge Graph (DKG) is modelled as a heterogeneous dynamic graph, represented by timestamped edge lists. Among the various formats for representing DKGs, Turtle (TTL) is one of the most commonly used. TTL allows the serialisation of nodes, edges, their attributes and timestamps, making it a versatile format for representing entities and relationships in a multi-agent system (MSA). An example of a TTL-based DKG is given in Eq. 1, where $s$ is the subject node, $o$ is the object node, $p$ is the predicate (or relationship), and $t$ is the timestamp.
$\text{DKG}=\{(s,p,o,t)\mid t\in[t_{1},t_{2},...,T]\}$ (1)
In the problem studied in this paper, anomalous data is represented by its entity $s$, the time interval during which the anomaly occurs $(t_{start},t_{end})$, and the class of the anomaly $c$. Table 1 shows all the anomaly types $c$ in the dataset.
Table 1: Description of system faults and their types.
Fault Type Description
Node CPU Hog Static Consumes the CPU resources of the targeted node.
Node Drain Dynamic Forces rescheduling of the resources running on
the targeted node.
Node Memory Hog Static Consumes the memory resources of the targeted
node.
Pod Autoscaler Dynamic Changes the desired number of replicas for the
targeted workload.
Pod CPU Hog Static Consumes the CPU resources of the targeted
application container.
Pod Delete Dynamic Causes forced/graceful pod failure of a random
replica of the targeted workload
Pod Network Latency Static Latency introduced in the network communication
between different pods.
2.2 Anomaly Detection
Anomaly detection is a crucial task in various fields, including network security, fraud detection, and financial systems [25, 26]. This task focuses on identifying anomalous patterns that deviate from the typical observation of the data [15].
Traditional anomaly detection methods typically rely on statistical techniques and distance-based measures. However, these methods often struggle with high-dimensional data and complex dependencies [27]. To overcome these problems, common machine learning approaches include Support Vector Machine (SVM) [6], eXtreme Gradient Boosting (XGB) [8], and Isolation Forest (IF) [7]. SVM identifies the optimal hyperplane that maximises the distance between normal data points and the decision boundary, effectively distinguishing outliers. XGB is a gradient-boosting decision tree algorithm that is able to evaluate the importance of each feature and find those that are most effective in separating normal from abnormal data. Isolation Forest is an unsupervised method that detects anomalies by constructing decision trees to isolate observations, with outliers requiring fewer splits to isolate due to their rarity and distinctiveness.
With the advancement of software algorithms and hardware, deep learning methods have been increasingly popular for anomaly detection due to their ability to extract complex features [16]. Common architectures for time series include Recurrent Neural Networks (RNN) [9], Temporal Convolutional Networks (TCN) [12], and the Self-Attention mechanism (SA) [13]. RNNs, including Long Short-Term Memory (LSTM) [10] and Gated Recurrent Unit (GRU) [11], capture temporal dependencies, addressing gradient issues in long sequences. TCNs handle long-term dependencies through causal convolutions, while SA captures long-range dependencies by examining relationships across sequences. For graph anomaly detection, Graph Neural Networks (GNN) propagate information through graph structures to identify anomalous patterns [19, 18].
Despite advancements in anomaly detection algorithms, significant challenges persist in dynamic knowledge graphs. Traditional machine learning methods struggle to capture the complex temporal and structural dependencies, while deep learning models require large amounts of labelled data, which is often scarce. Although techniques like negative sampling [19, 18, 20] can generate anomalous edges in dynamic graphs without attributed edges, this approach is infeasible for dynamic knowledge graphs with attributed edges. Capturing both temporal and structural dependencies remains key to tackling this challenge.
3 Methodology
This section introduces three representations of dynamic knowledge graphs: sequential data, one-hop, and two-hop graph structures, each adding increasing structural complexity. We then describe the machine learning and deep learning methods applied to these representations, followed by proposed ensemble learning schemes. Figures 1 and 2 provide an overview of the pipeline, including the data structure, the three levels of the data processing, and the final output.
Figure 1: Left: The ontology overview of the relationships between Kubernetes concepts of the ADDKG dataset. Right: The simplified hierarchical structure of the ADDKG dataset. The container information is aggregated as the attributes of Pods. The Connections serve as attributes of the edges between Pods and Services.
Figure 2: This figure presents the methodology used in this study, with three levels of processing: sequential data, one-hop, and two-hop graph structures. Each level introduces increasing structural complexity, showing how these representations are fed into anomaly detection models.
3.1 DKG as Sequential Data
The time-evolving attributes of nodes can be obtained by extracting raw TTL data, without considering the structural information of the DKG. As shown in the "DKG as Sequential Data" part of Fig. 2, we merge all the sequential data of all the vertices in a DKG by the timestamp. This includes vertices of type $\mathcal{C}$ representing "Cluster", type $\mathcal{P}$ representing "Pod", type $\mathcal{S}$ representing "Service", and type $\mathcal{N}$ representing "Node". For each given node category $k$, the time-evolving attributes can be represented as a tensor $\mathbf{\tilde{X}}_{k}\in\mathbb{R}^{N_{k}\times T\times d_{k}}$, where $N_{k}$ indicated the number of instances of the category $k$, $T$ indicates the number of snapshots, and $d_{k}$ indicates the number of features. A sliding window then captures temporal dependencies by concatenating the historical information of the past $\tau$ time steps, as shown in Eq. 2.
$\mathbf{f}(t_{j})=[\mathbf{x}(t_{j}-\tau+1)\,|\,\mathbf{x}(t_{j}-\tau+2)\,|\,\cdots\,|\,\mathbf{x}(t_{j})]$ (2)
Therefore, the sequential data representation of a given category $k$ in the DKG can be expressed by a matrix $\mathbf{X}\in\mathbb{R}^{T\times D_{k}}$, where $D_{k}=N_{k}\times d_{k}$, as shown in Eq. 3, where $\mathbf{f}(t_{j})$ is the aggregated attribute vector at timestamp $t_{j}$.
$\mathbf{X}=\begin{bmatrix}\mathbf{f}(t_{1})\\ \mathbf{f}(t_{2})\\ \vdots\\ \mathbf{f}(t_{T})\end{bmatrix}=\begin{bmatrix}\mathbf{x}(t_{1}-\tau+1)&\mathbf{x}(t_{1}-\tau+2)&\cdots&\mathbf{x}(t_{1})\\ \mathbf{x}(t_{2}-\tau+1)&\mathbf{x}(t_{2}-\tau+2)&\cdots&\mathbf{x}(t_{2})\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{x}(t_{T}-\tau+1)&\mathbf{x}(t_{m}-\tau+2)&\cdots&\mathbf{x}(t_{m})\end{bmatrix}$ (3)
Once the matrix representation of the attributes $\mathbf{X}$ is obtained, additional information can be computed. For example, the difference captures the trend of the time series, while the variance reflects the stability of the data fluctuations.
Based on this data representation, the temporal information across multiple time steps is aggregated into a single vector. Consequently, machine learning algorithms such as SVM, XGB, and IF, as well as deep learning models like MLP, TCN, and RNN, can use this input to capture information across multiple time steps.
3.2 One-Hop Graph Structure
To explore the structural information of the knowledge graph, we propose to use one-hop neighbour aggregation to capture the relationships between different entities. Given that the relationships in the DKG of an MSA are mainly hierarchical, i.e., a cluster contains multiple nodes, a node contains multiple pods, and a pod is connected to services through multiple connections (see Fig. 1), it is possible to aggregate information from leaf nodes to root nodes to obtain the structural-temporal features of each node at each timestamp, as shown the aggregate method in the "One-Hop Chart Structure" section of Fig. 2.
$S(t)=\bigcup_{i=1}^{n_{S}}\left(S_{i}(t),\bigcup_{j\in C(S_{i})}\left(C_{j}(t),\bigcup_{k\in P(C_{j})}\left(P_{k}(t),\bigcup_{l\in N(P_{k})}N_{l}(t)\right)\right)\right)$ (4)
For instance, to encode and predict data for the category Service, information related to the Service can be aggregated. Specifically, for all service instances $\mathcal{S}(t)$ at timestamp $t$, the dynamic information from related Nodes, Pods, Containers and Connections can be aggregated to form the attributes of the Service, as shown in Eq. 4, where $n_{S}$ represents the number of Service instances, $S_{i}$ represents the $i$-th instance of the Service category, $C_{j}(S_{i})$ represents the elements in the set of Connection instances related to $S_{i}$, $P_{k}(C_{j})$ represents the elements in the set of Pod instances related to $C_{j}$, and $N_{l}(P_{k})$ represents the elements in the set of Node instances related to $P_{k}$. In this manner, multiple one-hop graph structures are used for message propagation, and these structural-temporal features provide the model with a richer context.
3.3 Two-Hop Graph Structure
Some graph machine learning algorithms have pointed out that a larger convolutional field, or aggregation of multi-hop neighbours, can optimise the model. Inspired by the potential of aggregating second-hop neighbours to provide richer information, we observed that in the tree-like graph of the MSA problem, considering only the current level and its root forms a bipartite graph. For nodes of category $k$, the two-hop neighbours remain nodes of the same category. Recognising potential interdependencies between Service instances at each timestamp, we propose two-hop neighbour aggregation. This can be intuitively achieved by concatenating vertex features, as shown in Eq. 5 where $N_{\mathcal{S}}$ represents the number of Service instances and $S_{i}(t)$ represents the data for the $i$-th Service instance at timestamp $t$.
$\mathbf{X}_{\text{service}}(t)=[\mathbf{S_{1}}(t),\mathbf{S_{2}}(t),\ldots,\mathbf{S_{\mathit{N_{S}}}}(t)];$ (5)
Another method of capturing two-hop neighbours is using the Self-Attention mechanism (SA) to capture the correlations among vertices. SA works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others. This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data. See the "two hop graph structure" part in Fig. 2.
3.4 Ensemble Learning
Ensemble learning is a prevalent methodology in machine learning that reduce variance and enhance model robustness by combining multiple models. The effectiveness of ensemble learning depends on the integration of diverse, high-performing models [14]. It allows multiple models to make predictions on the same problem and vote on the final predictions. Voting can be hard or soft voting, with the mechanism such as unanimous and majority voting, which will be introduced in this subsection.
Figure 3: The ensemble learning framework used in this competition. The best-performing models: XGBoost (XGB), Support Vector Machine (SVM), and Self-Attention (SA) are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly. Optionally, a two-stage model combination can be applied, where the unsupervised Isolation Forest further refines the anomalies detected in $\hat{y_{1}}$ to reduce false positives.
In hard voting, each classifier votes for a class label, and the final prediction is determined by the majority of these classifiers. In soft voting, each classifier provides a probability for each class, and the final prediction is based on the highest average probability across all classifiers.
As two common voting mechanisms combined with hard voting, unanimous voting classifies an instance as an anomaly only if all the models agree. Majority voting, on the other hand, classifies an instance as an anomaly if the majority of models agree.
Intuitively, unanimous voting is more conservative and reduces false positives for anomaly detection, but may also miss some true anomalies. The majority voting method, in contrast, has some flexibility and can improve the detection of true anomalies, but it also preserves the false alarms. Both methods are tested on multiple sets of models and it are discussed in the next section.
4 Experiment
4.1 Experimental Setup
All experiments were conducted on the ISWC 2024 ADDKG dataset, containing five days of labelled data. Three days were used for training and two for testing, with data sampled at 15 second intervals. The dataset has a significant imbalance, with anomalies representing only 4% of the total. To tackle this, class weights were incorporated into the loss functions to improve the models’ ability of anomaly detection.
As outlined in the methodology section, we tested three levels of representation. For the first representation, sequential modelling is used, where the training set is randomly divided into 80% for training and 20% for validation. For the other two representations, a more refined data splitting strategy is used. While the 80-20 split is still used for the normal class, anomalies are treated as continuous anomalous events to preserve their integrity. This ensures that the models learn from complete sequences of anomalies. For each anomaly type, one complete event is randomly selected for validation, and the remaining events are used for training.
In the paper, we focus on binary anomaly detection, aiming to distinguish between normal and anomalous data without delving into the specific types of anomalies. The predictions are evaluated by F1 score, precision, and recall.
4.2 Performance of Models
This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, one-hop, and two-hop graph representations. Table 2 summarises the performance of individual models, while table 3 focuses on the results of different ensemble methods.
The first part of the experiments evaluates the performance of each model under different data representations, as shown in Table 2. The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 score.
Table 2: Performance of each model across different data representations. The best performance is highlighted in bold.
Model Dataset Phase F1 Score Precision Recall
Baseline - 0.06759 0.03498 1.00000
MLP D1 0.09390 0.05102 0.58824
XGB D2 0.21622 0.20000 0.23529
SVM D2 0.16185 0.08974 0.82353
TCN D1 0.09816 0.05479 0.47059
D2 0.07792 0.04206 0.52941
LSTM D1 0.06759 0.03498 1.00000
D2 0.10046 0.05446 0.64706
D3 0.16923 0.09735 0.64706
GRU D1 0.06759 0.03498 1.00000
D2 0.07362 0.04110 0.35294
D3 0.16250 0.09091 0.76471
Self-Attention D1 0.07203 0.03736 1.00000
D2 0.09028 0.04797 0.76471
D3 0.22018 0.13043 0.70588
The second part of the experiments evaluates the performance of different ensemble methods, as shown in table 3. We tested combinations of XGB, SVM, Self-Attention, and Isolation Forest, along with different voting mechanisms: soft, hard-unanimous, and hard-majority. Table 3 shows that the best performance is achieved with the combination of SVM, XGB and SA in hard-unanimous voting.
Table 3: Performance of different ensemble methods.The best performance is highlighted in bold.
Ensemble Model Method Dataset Phase F1 Score Precision Recall
Baseline - - 0.06759 0.03498 1.00000
XGB + IF soft D2 0.34286 0.33333 0.35294
SVM + IF hard,unanimous D2 0.19802 0.11905 0.58824
XGB + SVM + IF soft D2 0.38596 0.275 0.64706
XGB + SVM + SA + IF hard,unanimous D2+D3 0.36364 0.2963 0.47059
XGB + SVM + SA hard,unanimous D2+D3 0.51429 0.5 0.52941
Figure 4: The figure shows the F1 scores of the models across the three dataset levels, as well as the F1 scores obtained after applying ensemble learning methods. The individual models include XGB, SVM, TCN, LSTM, GRU, and the Self-Attention Model. The models used in the ensemble learning approach are XGB, Isolation Forest, SVM, and the Self-Attention Model.
Fig. 4 shows that the one-hop representation, two-hop representation as well as the ensemble learning methods all significantly outperforms the baseline result  [^5] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
5 Conclusion
In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a microservices environment. Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, one-hop, and two-hop graph structures.
Our results show that incorporating graph structures significantly enhances model performance. The two-hop graph structure, which includes service interactions, delivered the best results across most metrics, emphasising the importance of capturing complex interdependencies. Deep learning models, especially the Self-Attention model, performed particularly well in leveraging these structures.
Furthermore, the ensemble methods significantly boosted performance by combining the strengths of the top individual models. Unanimous voting, which requires all models to agree on an anomaly, outperformed majority voting in terms of precision, matching the competition’s emphasis on precision.
Future work should focus on several improvements. First, improving model scalability to handle larger, more complex datasets is essential. Second, exploring advanced ensemble techniques such as stacking or blending could improve detection accuracy. Third, integrating real-time anomaly detection would benefit practical applications in dynamic environments. Additionally, our study was limited to binary classification; future research should improve the accuracy of detecting different types of anomalies. Finally, although we used graph structure information, the application of graph neural networks remains a promising direction for further investigation.
This study highlights the potential of combining graph-based data representations with advanced machine learning and deep learning techniques for anomaly detection on dynamic knowledge graphs. Refining these methods and exploring the suggested future directions can lead to more robust, accurate, and scalable solutions for complex anomaly detection in various domains.